# Balancing properties of tropical moduli maps

### v1

@@ 2.2.2. Moduli of parameterized tropical curves
We denote by $M_{g,n,\nabla}^{\mathrm{trop}}$ the moduli space of parameterized stable tropical curves of genus $g$, degree $\nabla$, and with $n$ contracted legs. We always assume that the first $n$ legs $l_{1},\dotsc,l_{n}$ are the contracted ones. The space $M_{g,n,\nabla}^{\mathrm{trop}}$ is a generalized polyhedral complex with integral structure in the sense of [ACP15]. It is stratified by subsets $M_{[\Theta]}$ indexed by combinatorial types $\Theta$ with the fixed invariants $g,\nabla,$ and $n$. If ${\mathbb{G}}$ is the underlying graph of a combinatorial type $\Theta$, then $M_{[\Theta]}=M_{\Theta}/\mathrm{Aut}(\Theta)$, where $M_{\Theta}$ is the interior of a polyhedron $\overline{M}_{\Theta}$ in ${\mathbb{R}}^{|E({\mathbb{G}})|}\times N_{\mathbb{R}}^{|V({\mathbb{G}})|}$ and parameterizes tropical curves $h\colon\Gamma\to N_{{\mathbb{R}}}$ of type $\Theta$. The $e$-coordinate for an edge $e\in E({\mathbb{G}})$ is the length $\ell(e)$ of $e$; the $v$-coordinate for a vertex $v\in V({\mathbb{G}})$ is the coordinate of $h(v)\in N_{\mathbb{R}}$.
We call the stratum $M_{\Theta}$ weightless and $3$-valent if so is the type $\Theta$, and we call it weightless and almost $3$-valent if the type $\Theta$ is weightless and $3$-valent except for a unique $4$-valent vertex. If $M_{\Theta}$ is weightless and almost $3$-valent it is contained in the closure of at most $3$ strata, all of which are weightless and $3$-valent. Furthermore, it has codimension one in each of them.
@@ 2.2.3. Families of parameterized tropical curves over polyhedral complexes
For any $W\in\mathcal{F}(\Lambda)$ and $q\in W^{\circ}$, set $\Gamma_{q}:=(\mathbb{G}_{W},\ell_{q})$, where $\ell_{q}\colon E({\mathbb{G}}_{W})\to{\mathbb{R}}_{\geq 0}$ is the function defined by $\ell_{q}(\gamma):=\ell_{W}(\gamma,q)$. Denote by $h_{q}\colon\Gamma_{q}\rightarrow N_{\mathbb{R}}$ the unique piecewise affine map for which $h_{q}(u)=h_{W}(u,q)$ for all $u\in V({\mathbb{G}}_{W})$ and $\left(\frac{\partial h}{\partial\vec{l}_{i}}\right)=\overline{\nabla}$.
@@
Let $\nabla$ be the degree associated to the extended degree $\overline{\nabla}$ by removing the zero slopes. Any family of parameterized tropical curves $h\colon\Gamma_{\Lambda}\to N_{\mathbb{R}}$ induces a piecewise integral affine map $\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$ by sending $q\in\Lambda$ to the point parameterizing the isomorphism class of the fiber $h_{q}\colon\Gamma_{q}\to N_{\mathbb{R}}$. Furthermore, $\alpha$ lifts to an integral affine map from the interior of each face of $\Lambda$ to the corresponding stratum $M_{\Theta}$. The following definition is a generalization of [CHT23, Definition 3.4].
@@
    Suppose $\alpha({\rm Star}(W^{\circ}))\nsubseteq M_{[\Theta]}$. We say that $\alpha$ is locally combinatorially surjective at $W$ if for any combinatorial type $\Theta^{\prime}$ with an inclusion $\overline{M}_{\Theta}\hookrightarrow\overline{M}_{\Theta^{\prime}}$, we have $\alpha({\rm Star}(W^{\circ}))\cap M_{[\Theta^{\prime}]}\neq\emptyset$.
@@ 2.3. Families of curves
By a family of curves, we mean a flat, projective morphism of finite presentation and relative dimension one. By a collection of marked points on a family of curves, we mean a collection of disjoint sections contained in the smooth locus of the family. A family of curves with marked points is prestable if its geometric fibers have at-worst-nodal singularities. It is called (semi-)stable if so are its geometric fibers. A prestable curve with marked points defined over a field is called split if the irreducible components of its normalization are geometrically irreducible and smooth and the preimages of the nodes in the normalization are defined over the ground field. A family of prestable curves with marked points is called split if all of its fibers are so; cf. [dJ96, § 2.22]. If $U\subset Z$ is open, and $({\mathcal{C}},\sigma_{\bullet})$ is a family of curves with marked points over $U$, then by a model of $({\mathcal{C}},\sigma_{\bullet})$ over $Z$ we mean a family of curves with marked points over $Z$, whose restriction to $U$ is $({\mathcal{C}},\sigma_{\bullet})$; and similarly, for prestable and semi-stable models.
@@ 2.4. Toric varieties and parameterized curves
For a toric variety $X$, we denote the lattice of characters by $M$, of cocharacters by $N$, and the monomial functions by $x^{m}$ for $m\in M$. A parameterized curve in a toric variety $X$ is a smooth projective curve with marked points $(C,\sigma_{\bullet})$ together with a map $f\colon C\to X$ such that $f(C)$ does not intersect orbits of codimension greater than one, and the image of $C\setminus\left(\bigcup_{i}\sigma_{i}\right)$ under $f$ is contained in the dense orbit of $X$.
A family of parameterized curves $f\colon{\mathcal{C}}\rightarrow X$ over $K$ consists of the following data:
@@
    a map $f\colon{\mathcal{C}}\rightarrow X$, such that for any geometric point $p\in B$ the restriction ${\mathcal{C}}_{p}\to X$ is a parameterized curve.
@@ 2.5. Tropicalization of parameterized curves
Let $f\colon C\to X$ be a parameterized curve, and $C^{0}\to{\rm Spec}(K^{0})$ a prestable model. Denote by $\widetilde{C}$ the fiber of $C^{0}$ over the closed point of ${\rm Spec}(K^{0})$. The tropicalization $\mathrm{trop}(C)$ of $C$ with respect to the model $C^{0}$ is the tropical curve $\Gamma=(\mathbb{G},\ell)$ defined as follows: The underlying graph ${\mathbb{G}}$ is the dual graph of the central fiber $\widetilde{C}$, i.e., the vertices of $\mathbb{G}$ correspond to irreducible components of $\widetilde{C}$, the edges – to nodes, the legs – to marked points, and the natural incidence relation holds. For a vertex $v$ of $\mathbb{G}$, its weight is defined to be the geometric genus of the corresponding component $\widetilde{C}_{v}$ of the reduction $\widetilde{C}$. As for the length function, if $e\in E({\mathbb{G}})$ is the edge corresponding to a node $z\in\widetilde{C}$, then $\ell(e)$ is defined to be the valuation of $\lambda$, where $\lambda\in K^{00}$ is such that étale locally at $z$, the total space of $C^{0}$ is given by $xy=\lambda$. Although $\lambda$ depends on the étale neighborhood, its valuation does not, and hence the length function is well-defined. Finally, notice that the order on the set of marked points induces an order on the set of legs of $\Gamma$.
@@
Next, we explain how to construct the parameterization $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$. Let $\widetilde{C}_{v}$ be an irreducible component of $\widetilde{C}$. Then, for any $m\in M$, the pullback $f^{*}(x^{m})$ of the monomial $x^{m}$ is a non-zero rational function on $C^{0}$ since the preimage of the dense orbit is dense in $C$. Thus, there is a $\lambda_{m}\in K^{\times}$, unique up to an element invertible in $K^{0}$, such that $\lambda_{m}f^{*}(x^{m})$ is an invertible function at the generic point of $\widetilde{C}_{v}$. The function $h(v)$, associating to $m\in M$ the valuation $\nu(\lambda_{m})$, is clearly linear, and hence $h(v)\in N_{\mathbb{R}}$. The parameterization $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$ is defined to be the unique piecewise integral affine function with values $h(v)$ at the vertices of $\mathrm{trop}(C)$, whose slopes along the legs satisfy the following: for any leg $l$ and $m\in M$ we have $\frac{\partial h}{\partial\vec{l}}(m)=-\mathrm{ord}_{\sigma_{i}}f^{*}(x^{m})$, where $\sigma_{i}$ is the marked point corresponding to $l$. Then $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$ is a parameterized tropical curve, by [Tyo12, Lemma 2.23]. It is called the tropicalization of $f\colon C\to X$ with respect to the model $C^{0}$.
@@
Plainly, the tropical curve $\mathrm{trop}(C)$ is independent of the parameterization and depends only on $C^{0}$. If the family $C\to X$ is stable and $C^{0}$ is the stable model, then the corresponding tropicalization is called simply the tropicalization of $C$ (resp. of $f\colon C\to X$).
@@
Let $l$ be the leg corresponding to a marked point $\sigma_{i}$. If $\sigma_{i}$ is mapped to the boundary divisor $D$, then, by definition, $f(\sigma_{i})$ belongs to a unique component of $D$, and in particular, to the smooth locus of $D$. Thus, the direction of the slope $\frac{\partial h}{\partial\vec{l}}$ is completely determined by the irreducible component of $D$ containing the image of $\sigma_{i}$. Furthermore, the integral length of $\frac{\partial h}{\partial\vec{l}}$ is the multiplicity of $\sigma_{i}$ in $f^{*}D$. Note also that $\frac{\partial h}{\partial\vec{l}}=0$ if and only if $f(\sigma_{i})$ is contained in the dense orbit of $X$.
@@ 2.6. Strictly semistable pairs and their tropicalization
To tropicalize families, we will first tropicalize the base of the family. To do so, we shall use higher-dimensional tropicalizations. Different versions of such tropicalizations have appeared in the literature; see, for example, [MN15, ACGS20, CCUW20, Ran22, CGM22]. Here we will follow the construction by Gubler, Rabinoff, and Werner [GRW16]. Let us recall their construction.
@@
$D\coloneqq\widetilde{B}\cup(\bigcup_{k}H_{k})\subset B^{0}$ (2.2)
@@
For $U$ and $\Delta_{S}$ as above, consider the locus $U_{S}\subset U\setminus D$ of $K$-points $p$, whose specialization belongs to the union of strata $\cup_{S\leq T}T$. There is a natural tropicalization map $\tau_{U}\colon U_{S}\rightarrow\Delta_{S}$ given by
@@
The polyhedral complex $S(B^{0},H^{0}):=\bigcup_{S\in\mathrm{Str}(B^{0},H^{0})}\Delta_{S}$ is called the skeleton of the strictly semistable pair $(B^{0},H^{0})$.
@@ 3. Main theorem
In this section, we fix a pair of dual lattices $M$ and $N$, and a toric variety $X$ with lattice of monomials $M$. Let ¯$\overline{{\mathcal{M}}}_{g,n}$ be the algebraic stack over $\mathbb{Z}$ classifying stable $n$-pointed curves of genus $g$. Then ¯$\overline{{\mathcal{M}}}_{g,n}$ admits a finite surjective morphism from a projective scheme $\overline{M}$, over which there is a universal family of curves, cf. [dJ96, §2.24].
@@
Let $(B^{0},H^{0})$ be a strictly semistable pair over $K^{0}$ and $({\mathcal{C}}^{0}\rightarrow B^{0},\sigma^{0}_{\bullet})$ a split family of stable marked curves over $B^{0}$. Let $B$, $H$ and ${\mathcal{C}}$ be the generic fibers of $B^{0}$, $H^{0}$, and ${\mathcal{C}}^{0}$, respectively. Suppose ${\mathcal{C}}^{0}$ is obtained as a pullback of the universal family over $\overline{M}$ and is smooth over $B^{\prime}:=B\backslash H$. Suppose furthermore that $f\colon{\mathcal{C}}|_{B^{\prime}}\rightarrow X$ is a family of parameterized curves over $B^{\prime}$. Then there exists a family of parameterized tropical curves $h\colon\Gamma_{\Lambda}\to N_{\mathbb{R}}$ over $\Lambda:=S(B^{0},H^{0})$ such that
@@
    For any face $W\in{\mathcal{F}}(\Lambda)$ of codimension 1 (resp. of codimension at least 2) such that the tropical curves over the interior of $W$ are weightless and $3$-valent except for at most one $4$-valent vertex, the induced map $\alpha\colon\Lambda\to M^{\mathrm{trop}}_{g,n,\nabla}$ is either harmonic (resp. quasi-harmonic) or locally combinatorially surjective along $W$.
The following corollary is useful in applications, see, e.g., [CHT24].
@@
By [Sta20, Tag 0C2S], $B$ admits a flat projective model $B^{0}$ over $K^{0}$. Let $\overline{M}^{0}$ be the trivial model of $\overline{M}$ over $K^{0}$. After replacing $B^{0}$ with the schematic image of the graph $B\to\overline{M}^{0}\times_{{\rm Spec}(K^{0})}B^{0}$, which is also flat over $K^{0}$ by loc. cit., we may assume that the family of marked curves extends to $B^{0}$. By [dJ96, Lemma 5.3], there exist a projective alteration $\varphi_{1}\colon B_{1}^{0}\to B^{0}$ such that the pullback of ${\mathcal{C}}$ to $B_{1}^{0}$ is split over $B_{1}^{0}$. Replacing $B_{1}^{0}$ with an irreducible component that dominates $B^{0}$ if necessary, we may assume that $B_{1}^{0}$ is integral. Since the loci $\Sigma$ associated to $B^{\prime}$ and to $\varphi_{1}^{-1}(B^{\prime})$ coincide, we may assume that $B_{1}^{0}=B^{0}$ and hence the family ${\mathcal{C}}$ is split over $B^{0}$. Finally, by [dJ96, Theorem 6.5], there exists an alteration $\varphi_{2}\colon B_{2}^{0}\to B^{0}$ such that $B_{2}^{0}$ is projective over $K^{0}$, the generic fiber of $B_{2}^{0}\to{\rm Spec}(K^{0})$ is irreducible, and the pair $\left(B_{2}^{0},\overline{\varphi_{2}^{-1}(H)}_{\rm red}\right)$ is strictly semistable. Here $\overline{\varphi_{2}^{-1}(H)}_{\rm red}$ is the closure of $\varphi_{2}^{-1}(H)_{\rm red}$ in $B_{2}^{0}$. As before, we may assume that $B_{2}^{0}=B^{0}$.
@@
(2) Set $k:=\dim M_{[\Theta]}$, and assume to the contrary that $\overline{M}_{[\Theta]}\nsubseteq\alpha(\Lambda)$. By (1), $\alpha(\Lambda)\subseteq M^{\mathrm{trop}}_{g,n,\nabla}$ is closed, and therefore $M_{[\Theta]}\nsubseteq\alpha(\Lambda)$. Furthermore, there is a face $W\in{\mathcal{F}}(\Lambda)$ such that $\alpha(W^{\circ})$ has dimension $k-1$, $\alpha\left({\rm Star}(W^{\circ})\right)$ has dimension $k$, and $\alpha(W^{\circ})$ belongs to the relative boundary of $\alpha(\Lambda)$ in $M_{[\Theta]}$.
After replacing $W$ with one of its faces, we may assume that $\dim(W)=k-1$. Indeed, since $\dim(\alpha(W^{\circ}))=k-1$, the dimension of $W$ is at least $k-1$. If it is greater than $k-1$, then the kernel of the differential $d\alpha$ along $W$ is non-trivial. But $\Lambda$ is the skeleton of a strictly semistable pair, and therefore $W$ contains no lines. Thus, $\alpha(W)=\alpha(\partial W)$, and by induction, we can find a face of $W$ of dimension $k-1$ satisfying the requirements.
@@
The above results, while applicable to the more general case of weightless and (almost) $3$-valent strata, are conceived with a view towards their application in [CHT24]. There we will use simple walls and nice strata of [CHT23], which are weightless and (almost) $3$-valent strata, which in addition are regular, i.e., of the expected dimension. Two weightless and (almost) $3$-valent, but not regular, strata were studied already in [CHT22]:

### v2

@@ 2.2.2. Moduli of parameterized tropical curves
Fixing the genus $g$, degree $\nabla$, and number $n$ of contracted legs, we denote by $M_{g,n,\nabla}^{\mathrm{trop}}$ the moduli space of parameterized stable tropical curves with these invariants. We will always assume that the first $n$ legs $l_{1},\dotsc,l_{n}$ are the contracted ones. Then the space $M_{g,n,\nabla}^{\mathrm{trop}}$ is a generalized polyhedral cone complex with integral structure in the sense of [ACP15], whose strata $M_{[\Theta]}$ are indexed by combinatorial types $\Theta$ with the fixed invariants. If $\mathbb{G}$} is the underlying graph of the combinatorial type $\Theta$, then $M_{[\Theta]}=M_{\Theta}/\mathrm{Aut}(\Theta)$, where $\mathrm{Aut}(\Theta)$ denotes the automorphisms of the combinatorial type. Here $M_{\Theta}$ is the interior of a polyhedron $\overline{M}_{\Theta}$ in ${\mathbb{R}}^{|E({\mathbb{G}})|}\times N_{\mathbb{R}}^{|V({\mathbb{G}})|}$ and it parameterizes tropical curves $h\colon\Gamma\to N_{{\mathbb{R}}}$ of type $\Theta$: the $e$-coordinate for an edge $e\in E({\mathbb{G}})$ encodes the length $\ell(e)$ of $e$; the $v$-coordinate for a vertex $v\in V({\mathbb{G}})$ encodes the coordinate of $h(v)\in N_{\mathbb{R}}$.
We call the stratum $M_{\Theta}$ weightless and $3$-valent if so is the combinatorial type $\Theta$, and we call it weightless and almost $3$-valent if $\Theta$ is weightless and $3$-valent except for a unique $4$-valent vertex. If $M_{\Theta}$ is weightless and almost $3$-valent it is contained in the closure of at most $3$ strata, all of which are weightless and $3$-valent. Furthermore, it has codimension one in each of them.
@@ 2.2.3. Families of parameterized tropical curves over polyhedral complexes
For any $W\in\mathcal{F}(\Lambda)$ and any $q\in W^{\circ}$, set $\Gamma_{q}:=(\mathbb{G}_{W},\ell_{q})$, where $\ell_{q}\colon E({\mathbb{G}}_{W})\to{\mathbb{R}}_{\geq 0}$ is the function defined by $\ell_{q}(\gamma):=\ell_{W}(\gamma,q)$. Denote by $h_{q}\colon\Gamma_{q}\rightarrow N_{\mathbb{R}}$ the unique piecewise affine map for which $\left(\frac{\partial h}{\partial\vec{l}_{i}}\right)=\overline{\nabla}$$h_{q}(u)=h_{W}(u,q)$ for all $u\in V({\mathbb{G}}_{W})$ and $\left(\frac{\partial h}{\partial\vec{l}_{i}}\right)=\overline{\nabla}$.
@@
Remark 2.8.
The version of tropical families and tropicalization we use in the current paper is natural when working with families of parametrized curves over non-Archimedean fields as the tropicalization of the base embeds naturally as a skeleton in the Berkovich analytification of the base, and similarly for the fibers. There exists another version of tropicalization arising naturally when working in the logarithmic category, see, e.g., [ACGS20, CCUW20, R17, Ran22]. The logarithmic tropicalization is more general, but the base and the fibers belong to the category of (generalized) polyhedral cone complexes. The relation between the two is given by considering the cones over the faces of $\Lambda$ and the cones over the tropical curves in the family. Notice also that the parametrization maps we consider in this paper are particular cases of (stable) maps in loc. cit.
@@
Let $\nabla$ be the degree associated to the extended degree $\overline{\nabla}$ by removing the zero slopes. Any family of parameterized tropical curves $h\colon\Gamma_{\Lambda}\to N_{\mathbb{R}}$ induces a piecewise integral affine map $\alpha\colon\Lambda\to M_{g,n,\nabla}^{\mathrm{trop}}$ by sending $q\in\Lambda$ to the point parameterizing the isomorphism class of the stabilization of the fiber $h_{q}\colon\Gamma_{q}\to N_{\mathbb{R}}$. Furthermore, $\alpha$ lifts to an integral affine map from the interior of each face of $\Lambda$ to the corresponding stratum $M_{\Theta}$. The following definition is a generalization of [CHT23, Definition 3.4].
@@
    Suppose $\alpha({\rm Star}(W^{\circ}))\nsubseteq M_{[\Theta]}$. We say that $\alpha$ is locally combinatorially surjective at $W$ if for any $\overline{M}_{[\Theta]}\subseteq\overline{M}_{[\Theta^{\prime}]}$, we have $\alpha({\rm Star}(W^{\circ}))\cap M_{[\Theta^{\prime}]}\neq\emptyset$.
@@ 2.3. Families of curves
A family of curves is a flat, projective morphism of finite presentation and relative dimension one; cf. [CHT22, § 2.1]. In our setting it often comes with marked points, that is, an ordered collection of disjoint sections contained in the smooth locus of the family. A family of curves with marked points is prestable if its geometric fibers have at-worst-nodal singularities; and (semi-)stable if so are its geometric fibers. A prestable curve with marked points defined over a field is split if the irreducible components of its normalization are geometrically irreducible and smooth, and the preimages of the nodes in the normalization are defined over the ground field. A family of prestable curves with marked points is called split if all of its fibers are so; cf. [dJ96, § 2.22]. Let $U\subset Z$ be an open subset, and $({\mathcal{C}},\sigma_{\bullet})$ a family of curves with marked points over $U$. Then a model of $({\mathcal{C}},\sigma_{\bullet})$ over $Z$ is a family of curves with marked points over $Z$, whose restriction to $U$ is $({\mathcal{C}},\sigma_{\bullet})$.
@@ 2.4. Toric varieties and parameterized curves
Throughout the paper, $M$ and $N$ are the lattice of characters, respectively cocharacters, of a toric variety $X$. The monomial functions are denoted by $x^{m}$ for $m\in M$. A parameterized curve in the toric variety $X$ is a smooth projective curve with marked points $(C,\sigma_{\bullet})$ together with a map $f\colon C\to X$ such that $f(C)$ does not intersect torus orbits of codimension greater than one, and the image of $C\setminus\left(\bigcup_{i}\sigma_{i}\right)$ under $f$ is contained in the dense orbit of $X$.
A family of parameterized curves $f\colon{\mathcal{C}}\rightarrow X$ consists of the following data:
@@
    a map $f\colon{\mathcal{C}}\rightarrow X$, such that for any geometric point $p\in B$ the restriction ${\mathcal{C}}_{p}\to X$ is a parameterized curve in the above sense.
@@ 2.5. Tropicalization of parameterized curves
The canonical tropicalization of parameterized curves is well established; see e.g. [Tyo12, CHT23]. We recall the construction here for convenience. Let $f\colon C\to X$ be a parameterized curve, $C^{0}\to{\rm Spec}(K^{0})$ a prestable model, and $\widetilde{C}$ the fiber of $C^{0}$ over the closed point of ${\rm Spec}(K^{0})$. The tropicalization $\mathrm{trop}(C)$ of $C$ with respect to the model $C^{0}$ is a tropical curve and we need to specify its underlying weighted graph $\mathbb{G}$}, as well as a length function $\ell$. The former is the dual graph of the central fiber $\widetilde{C}$, i.e., the vertices of $\mathbb{G}$ correspond to irreducible components of $\widetilde{C}$, the edges to nodes, and the legs to marked points. Together with the natural incidence relations this gives a finite graph $\mathbb{G}$, and the order on the set of marked points gives the order on its set of legs. The weight of a vertex $v$ of $\mathbb{G}$ is defined to be the geometric genus of the corresponding component $\widetilde{C}_{v}$ of the reduction $\widetilde{C}$. Finally, in order to define the length function on an edge $e\in E({\mathbb{G}})$, let $z\in\widetilde{C}$ be the node corresponding to $e$. Then étale locally at $z$, the total space of $C^{0}$ is given by $xy=\lambda$ with $\lambda\in K^{00}$. We set $\ell(e)=\nu(\lambda)$. Although $\lambda$ depends on the étale neighborhood, its valuation does not, and hence the length function is well-defined.
@@
Next, we explain how to construct the parameterization $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$. To this end, observe that for any $m\in M$ the pullback $f^{*}(x^{m})$ of the monomial $x^{m}$ gives a non-zero rational function on $C^{0}$ because the preimage of the dense orbit of $X$ is dense in $C$. Thus, for any irreducible component $\widetilde{C}_{v}$ of $\widetilde{C}$, there is a $\lambda_{m}\in K^{\times}$ such that $\lambda_{m}f^{*}(x^{m})$ is an invertible function at the generic point of $\widetilde{C}_{v}$. Here $\lambda_{m}$ is unique up to an element of valuation 0. The function $h(v)$, which associates to $m\in M$ the valuation $\nu(\lambda_{m})$, is clearly linear, and hence $h(v)\in N_{\mathbb{R}}$. The parameterization $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$ then is the unique piecewise integral affine function with values $h(v)$ at the vertices of $\mathrm{trop}(C)$, whose slopes along the legs are given as follows: for any leg $l$ and $m\in M$ we set $\frac{\partial h}{\partial\vec{l}}(m)=-\mathrm{ord}_{\sigma_{i}}f^{*}(x^{m})$, where $\sigma_{i}$ is the marked point corresponding to $l$. By [Tyo12, Lemma 2.23], $h\colon\mathrm{trop}(C)\to N_{\mathbb{R}}$ then is a parameterized tropical curve, that is, defines a harmonic map. It is called the tropicalization of $f\colon C\to X$ with respect to the model $C^{0}$.
@@
Clearly, the tropical curve $\mathrm{trop}(C)$ is independent of the parameterization and depends only on $C^{0}$. If the family $C\to X$ is stable and $C^{0}$ is the stable model, then the corresponding tropicalization is called simply the tropicalization of $C$ (resp. of $f\colon C\to X$).
@@
Let $l$ be a leg corresponding to a marked point $\sigma_{i}$. If $f$ maps $\sigma_{i}$ to the boundary divisor $D$ of $X$, then, by definition of parameterized curves, $f(\sigma_{i})$ belongs to a unique component of $D$, and in particular, to the smooth locus of $D$. It follows that the direction of the slope $\frac{\partial h}{\partial\vec{l}}$ is determined by the irreducible component of $D$ that contains the image of $\sigma_{i}$. Furthermore, the integral length of $\frac{\partial h}{\partial\vec{l}}$ equals the multiplicity of $\sigma_{i}$ in $f^{*}D$. Note also that $\frac{\partial h}{\partial\vec{l}}=0$ if and only if $f(\sigma_{i})$ is contained in the dense orbit of $X$.
@@ 2.6. Strictly semistable pairs and their tropicalization
To tropicalize families, we will first tropicalize the base of the family. To do so, we shall use higher-dimensional tropicalizations. As mentioned in Remark 2.8, different versions of such tropicalizations have appeared in the literature; see, for example, [MN15, ACGS20, CCUW20, Ran22, CGM22]. Here we will follow the construction by Gubler, Rabinoff, and Werner [GRW16]. Let us recall their construction.
@@
$D\coloneqq\widetilde{B}\cup\left(\bigcup_{k}H_{k}\right)\subset B^{0}$ (2.2)
@@
For $U$ and $\Delta_{S}$ as above, consider the locus $U_{S}\subset U\setminus D$ of $K$-points $p$, whose specialization belongs to the union of strata $\bigcup_{S\leq T}T$. There is a natural tropicalization map $\tau_{U}\colon U_{S}\rightarrow\Delta_{S}$ given by
@@
([GRW16, § 4.6]) The polyhedral complex $S(B^{0},H^{0}):=\bigcup_{S\in\mathrm{Str}(B^{0},H^{0})}\Delta_{S}$ is called the skeleton of the strictly semistable pair $(B^{0},H^{0})$.
Remark 2.14.
The choice of the term “skeleton” is due to the fact that the polyhedral complex $S(B^{0},H^{0})$ embeds naturally into the Berkovich analytification $B^{\rm an}$ of $B$, and its image is a strong deformation retract of $B^{\rm an}$.
@@ 3. Main theorem
In this section, we fix a pair of dual lattices $M$ and $N$, and a toric variety $X$ with lattice of monomials $M$. Let $\overline{{\mathcal{M}}}_{g,n}$ be the algebraic stack over $\mathbb{Z}$ classifying stable $n$-pointed curves of genus $g$. Then $\overline{{\mathcal{M}}}_{g,n}$ admits a finite surjective morphism from a projective scheme $\overline{M}$, over which there is a universal family of curves, cf. [dJ96, §2.24].
@@
[…]
@@
    For any face $W\in{\mathcal{F}}(\Lambda)$ of codimension one (resp. of codimension at least two) such that the tropical curves over the interior of $W$ are weightless and $3$-valent except for at most one $4$-valent vertex, the induced map $\alpha\colon\Lambda\to M^{\mathrm{trop}}_{g,n,\nabla}$ is either harmonic (resp. quasi-harmonic) or locally combinatorially surjective along $W$.
Remark 3.2.
Property (1) can be considered as the “defining property” of the tropicalization of a family. Indeed, any natural tropicalization procedure should commute with point-wise tropicalization. In fact, in the case of families of curves, the existence of tropicalizations satisfying property (1) follows from the (more general) logarithmic tropicalization construction of Cavalieri, Chan, Ulirsch, and Wise, [CCUW20, § 7.1-7.3]. Constructions of tropicalizations of families of stable maps satisfying (1) can also be found in the literature but under the assumption that the characteristic of the ground field is zero, see, e.g., [ACGS20] or [R17, Ran22].
The main new ingredient in Theorem 3.1 is the balancing properties of the induced map $\alpha$ described in (2). These balancing properties are crucial in some applications. For example, they can be used to prove realizability of parameterized curves in given families, and to prove irreducibility results for moduli of parameterized curves. In particular, the following corollary is used in the proof of irreducibility of Hurwitz schemes in any characteristic, [CHT25].
@@
By [Sta25, Tag 0C2S], $B$ admits a flat projective model $B^{0}$ over $K^{0}$. Let $\overline{M}^{0}$ be the trivial model of $\overline{M}$ over $K^{0}$. After replacing $B^{0}$ with the schematic image of the graph $B\to\overline{M}^{0}\times_{{\rm Spec}(K^{0})}B^{0}$, which is also flat over $K^{0}$ by loc. cit., we may assume that the family of marked curves extends to $B^{0}$. By [dJ96, Lemma 5.3], there exist a projective alteration $\varphi_{1}\colon B_{1}^{0}\to B^{0}$ such that the pullback of $\mathcal{C}$} to $B_{1}^{0}$ is split over $B_{1}^{0}$. Replacing $B_{1}^{0}$ with an irreducible component that dominates $B^{0}$ if necessary, we may assume that $B_{1}^{0}$ is integral. Since the loci $\Sigma$ associated to $B^{\prime}$ and to $\varphi_{1}^{-1}(B^{\prime})$ coincide, we may assume that $B_{1}^{0}=B^{0}$ and hence the family $\mathcal{C}$} is split over $B^{0}$. Finally, by [dJ96, Theorem 6.5], there exists an alteration $\varphi_{2}\colon B_{2}^{0}\to B^{0}$ such that $B_{2}^{0}$ is projective over $K^{0}$, the generic fiber of $B_{2}^{0}\to{\rm Spec}(K^{0})$ is irreducible, and the pair $\left(B_{2}^{0},\overline{\varphi_{2}^{-1}(H)}_{\rm red}\right)$ is strictly semistable. Here $\overline{\varphi_{2}^{-1}(H)}_{\rm red}$ is the closure of $\varphi_{2}^{-1}(H)_{\rm red}$ in $B_{2}^{0}$. As before, we may assume that $B_{2}^{0}=B^{0}$.
@@
(2) Set $k:=\dim M_{[\Theta]}$, and assume to the contrary that $\overline{M}_{[\Theta]}\nsubseteq\alpha(\Lambda)$. By (1), $\alpha(\Lambda)\subseteq M^{\mathrm{trop}}_{g,n,\nabla}$ is closed, and therefore $M_{[\Theta]}\nsubseteq\alpha(\Lambda)$. Furthermore, there is a face $W\in{\mathcal{F}}(\Lambda)$ such that $\alpha(W^{\circ})$ has dimension k1k-1, $\alpha\left({\rm Star}(W^{\circ})\right)$ has dimension $k$, and $\alpha(W^{\circ})$ belongs to the relative boundary of $\alpha(\Lambda)$ in $M_{[\Theta]}$.
After replacing $W$ with one of its faces, we may assume that $\dim(W)=k-1$. Indeed, since $\dim(\alpha(W^{\circ}))=k-1$, the dimension of $W$ is at least k1k-1. If it is greater than k1k-1, then the kernel of the differential $d\alpha$ along $W$ is non-trivial. But $\Lambda$ is the skeleton of a strictly semistable pair, and therefore $W$ contains no lines. Thus, $\alpha(W)=\alpha(\partial W)$, and by induction, we can find a face of $W$ of dimension k1k-1 satisfying the requirements.
@@
The above results, while applicable to the more general case of weightless and (almost) $3$-valent strata, are conceived with a view towards their application in [CHT25]. There we will use simple walls and nice strata of [CHT23], which are weightless and (almost) $3$-valent strata, which in addition are regular, i.e., of the expected dimension. Two weightless and (almost) $3$-valent, but not regular, strata were studied already in [CHT22]:
@@
Remark 3.5.
As already mentioned in Remark 3.2, the existence of tropicalizations of families of (parameterized) curves satisfying property (1) is known to the experts, see e.g., [ACGS20, CCUW20, R17, Ran22]. However, the results in loc. cit. are stated in a different language and some are proven under the assumption that the characteristic is zero. Therefore, we include a proof of the existence of tropicalizations satisfying (1) for the convenience of the reader.